# Computation of Greeks under rough Volterra stochastic volatility models using the Malliavin calculus approach

### v1

@@ 4.2 Mixed $\alpha$RFSV model
Moreover, we have
@@ 4.3 Rough Stein-Stein model
In relation with the existence and uniqueness of solution of this equation we refer the reader to Nualart and Ouknine (2002). Being the drift term derivable the Malliavin differentiability of the solution is straightforward.
@@
$\int_{0}^{T}G(t,T)\,\mathrm{d}t=\int_{0}^{T}V_{t}\,\mathrm{d}t+\rho\nu\int_{0}^{T}\left(\int_{0}^{s}D_{t}V_{s}\,\mathrm{d}t\right)(\,\mathrm{d}W_{s}-V_{s}\,\mathrm{d}s)$
@@
$\int_{0}^{T}\int_{0}^{T}D_{s}G(t,T)\,\mathrm{d}s\,\mathrm{d}t=\int_{0}^{T}\int_{0}^{T}D_{s\wedge t}V_{s\vee t}\,\mathrm{d}s\,\mathrm{d}t-\int_{0}^{T}\left(\int_{0}^{u}D_{s}V_{u}\,\mathrm{d}s\right)^{2}\,\mathrm{d}u.$
@@ 5 Conclusion
We used Malliavin calculus techniques to obtain formulas for computing Greeks under different rough Volterra stochastic volatility models. In particular we showed that each model is fully characterized by the quantity $G(t,T)$, defined in (20), and by its Malliavin derivative $D_{s}G(t,T)$ (see Lemma 3.2). Calculating the integrals (24) and (25) then leads us directly to all the Greeks formulas listed in Theorem 3.3 and Proposition 3.4. In particular we derived formulas for rough versions of the Stein-Stein, SABR and Bergomi models.
@@
The work of Josep Vives was partially supported by Spanish grant PID2020-118339GB-100 (2021-2024).
@@ Acknowledgements
Computational and storage resources were supplied by the project “e-Infrastruktura CZ” (e-INFRA LM2018140) supported by the Ministry of Education, Youth and Sports of the Czech Republic.
We thank professor David Nualart for some comments in relation with the reference Nualart and Ouknine (2002).
@@ A.2 Stein-Stein model
$=$ $V_{t}+\rho\nu\int_{t}^{T}$
$e^{-\kappa(s-t)}\,\mathrm{d}W_{s}-\rho\nu\int_{t}^{T}V_{s}\mathrm{e}^{-\kappa(s-t)}\,\mathrm{d}s$

### v2

@@ 4.2 Mixed $\alpha$RFSV model
In order to get all the ingredients needed to compute the Greeks for this model, one should repeat the computations for the $\alpha$RFSV model, add them and divide them by 2. Examples of how these computations work are, for example
@@ 4.3 Rough Stein-Stein model
In relation with the existence and uniqueness of solution of this equation we refer the reader to Nualart and Ouknine (2002). Being the drift term derivable, the Malliavin differentiability of the solution is straightforward.
@@
$\int_{0}^{T}G(u,T)\,\mathrm{d}u=\int_{0}^{T}V_{u}\,\mathrm{d}u+\rho\nu\int_{0}^{T}\left(\int_{0}^{s}D_{t}V_{u}\,\mathrm{d}u\right)(\,\mathrm{d}W_{s}-V_{s}\,\mathrm{d}s),$
$\int_{0}^{T}\int_{0}^{T}D_{s}G(u,T)\,\mathrm{d}u\,\mathrm{d}s=\int_{0}^{T}\int_{0}^{T}D_{s\wedge u}V_{s\vee u}\,\mathrm{d}s\,\mathrm{d}u-\int_{0}^{T}\left(\int_{0}^{u}D_{s}V_{u}\,\mathrm{d}s\right)^{2}\,\mathrm{d}u.$
@@
$\int_{0}^{T}\int_{0}^{T}\int_{0}^{T}D_{t}D_{s}G(u,t)\,\mathrm{d}u\,\mathrm{d}s\,\mathrm{d}t=0.$
In order to compute the Greeks $\mathcal{V}$ and $\mathcal{H}$, we need to compute the derivatives with respect to $V_{0}$ and $H$ of the quantities $V_{t}$, $\frac{-V_{t}^{2}}{2}$, $D_{t}V_{s}$ and $D_{t}\left(\frac{-V_{t}^{2}}{2}\right)$. On the one hand, notice that
$\partial_{V_{0}}V_{t}=1-\kappa\int_{0}^{t}\partial_{V_{0}}V_{u}\,\mathrm{d}u.$
Hence,
$a_{V_{0}}(V_{t})=\partial_{V_{0}}V_{t}=\mathrm{e}^{-\kappa t},\quad D_{s}a_{V_{0}}(V_{u})=0.$
In a similar fashion, we have that
$b_{V_{0}}(V_{t})=-\mathrm{e}^{-\kappa t}V_{t},\quad D_{t}b_{V_{0}}(V_{s})=-\mathrm{e}^{-\kappa s}\left(\rho\nu K_{H}(s,t)-\kappa\rho\nu\int_{t}^{s}K_{H}(u,t)\mathrm{e}^{-\kappa(s-u)}\,\mathrm{d}u\right).$
Concerning the derivatives with respect to $H$ we have
$\partial_{H}V_{t}=-\kappa\int_{0}^{t}\partial_{H}V_{u}\,\mathrm{d}u+\int_{0}^{t}\partial_{H}K_{H}(t,u)\,\mathrm{d}Z_{u}.$ (43)
Since
$\partial_{H}K_{H}(t,s)=\sqrt{2H}(t-s)^{H-1/2}\left(\frac{1}{2H}+\ln(t-s)\right)\in L^{2}([0,t])$
for all $t\in[0,T]$ we have that there exists a unique solution to equation (43). This implies that $a_{H}(V_{t})$ is the solution to (43) and $D_{t}a_{H}(V_{s})$ can be expressed as
$D_{t}a_{H}(V_{s})=\partial_{H}D_{t}V_{s}=\rho\nu\partial_{H}K_{H}(s,t)-\kappa\rho\nu\int_{t}^{s}\partial_{H}K_{H}(u,t)\mathrm{e}^{-\kappa(s-u)}\,\mathrm{d}u.$
Moreover, since $b_{H}(V_{t})=-V_{t}\partial_{H}V_{t}$, we have
$D_{t}b_{H}(V_{s})=$ $-D_{t}V_{s}\cdot\partial_{H}V_{s}-V_{t}\cdot\partial_{H}(D_{t}V_{s})$
$=$ $\partial_{H}V_{s}\left(\rho\nu K_{H}(s,t)-\kappa\rho\nu\int_{t}^{s}K_{H}(u,t)\mathrm{e}^{-\kappa(s-u)}\,\mathrm{d}u\right)$
-- $V_{s}\left(\rho\nu\partial_{H}K_{H}(s,t)-\kappa\rho\nu\int_{t}^{s}\partial_{H}K_{H}(u,t)\mathrm{e}^{-\kappa(s-u)}\,\mathrm{d}u\right).$
In conclusion, the ingredients needed to compute the approximate Greek formulas for the Stein-Stein model are the following:
  • $\mathcal{G}_{T}=\int_{0}^{T}V_{u}\,\mathrm{d}u+\rho\nu\int_{0}^{T}\left(\int_{0}^{s}D_{t}V_{u}\,\mathrm{d}u\right)(\,\mathrm{d}W_{s}-V_{s}\,\mathrm{d}s)$,
  • $\int_{0}^{T}D_{s}\mathcal{G}_{T}\,\mathrm{d}s=\int_{0}^{T}\int_{0}^{T}D_{s\wedge u}V_{s\vee u}\,\mathrm{d}s\,\mathrm{d}u-\int_{0}^{T}\left(\int_{0}^{u}D_{s}V_{u}\,\mathrm{d}s\right)^{2}\,\mathrm{d}u$,
  • $\int_{0}^{T}\int_{0}^{T}D_{t}D_{s}\mathcal{G}_{T}\,\mathrm{d}s\,\mathrm{d}t=0$,
  • $a_{V_{0}}(V_{t})=\mathrm{e}^{-\kappa t}$,
  • $b_{V_{0}}(V_{t})=-\mathrm{e}^{-\kappa t}V_{t}$,
  • $D_{t}a_{V_{0}}(V_{s})=0$,
  • $D_{t}b_{V_{0}}(V_{s})=-\mathrm{e}^{-\kappa s}\left(\rho\nu K_{H}(s,t)-\kappa\rho\nu\int_{t}^{s}K_{H}(u,t)\mathrm{e}^{-\kappa(s-u)}\,\mathrm{d}u\right)$.
  • $a_{H}(V_{t})=\partial_{H}V_{t}$, where $\partial_{H}V_{t}$ is the solution to equation (43).
  • $b_{H}(V_{t})=-V_{t}\partial_{H}V_{t}$.
  • $D_{t}a_{H}(V_{s})=\rho\nu\partial_{H}K_{H}(s,t)-\kappa\rho\nu\int_{t}^{s}\partial_{H}K_{H}(u,t)\mathrm{e}^{-\kappa(s-u)}\,\mathrm{d}u$.
  • $D_{t}b_{H}(V_{s})=$ $\partial_{H}V_{s}\left(\rho\nu K_{H}(s,t)-\kappa\rho\nu\int_{t}^{s}K_{H}(u,t)\mathrm{e}^{-\kappa(s-u)}\,\mathrm{d}u\right)$
    $-V_{s}\left(\rho\nu\partial_{H}K_{H}(s,t)-\kappa\rho\nu\int_{t}^{s}\partial_{H}K_{H}(u,t)\mathrm{e}^{-\kappa(s-u)}\,\mathrm{d}u\right).$
@@ 5 Conclusion
We used Malliavin calculus techniques to obtain formulas for computing Greeks under different rough Volterra stochastic volatility models. In particular, we showed that each model is fully characterized by the quantity $G(t,T)$, defined in (26), and by its Malliavin derivative $D_{s}G(t,T)$ (see Lemma 3.2). Calculating the integrals (30) and (31) then leads us directly to all the Greeks formulas listed in Theorem 3.3 and Proposition 4.1. In particular, we derived formulas for rough versions of the Stein-Stein, SABR and Bergomi models.
For some of the stochastic volatility models considered in the present paper, the value of the underlying price at $T$, $S_{T}$, does not belong to $L^{2}(\Omega)$. For example for rough Bergomi model, see Gassiat (2019). This justifies the extension of the integration by parts formula to the $L^{1}$ case in Theorem 2.8 and the use of an approximation of the unity in the proof of Theorem 3.3.
We derive the formulas for all five primary Greeks $\Delta$, $\Gamma$, $\varrho$, $\mathcal{V}$ and $\mathcal{H}$. We assumed that our price process $S$ followed (32) and the volatility or the variance $V$ followed (36). Explicitly, we calculate the Greeks for the $\alpha$RFSV model, that includes rough Bergomi and rough SABR models in some cases, the mixed $\alpha$RFSV model and the rough Stein and Stein model.
As numerical results, we calculated the Delta for $\alpha$RFSV model for European call and put. As we can see in Figure 1, Monte Carlo simulations converge satisfactorily. Moreover, for the first time we could present a dependence of Delta on the Hurst parameter $H$ in Figure 2.
@@
The work of Josep Vives was partially supported by Spanish grant PID2020-118339GB-100 (2021-2024). Author Òscar Burés is supported by the predoctoral program AGAUR-FI ajuts (2025 FI-1 00580)
@@ Acknowledgements
Computational and storage resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic.
We thank Professor David Nualart for some comments in relation to the reference Nualart and Ouknine (2002).
@@ A.2 Stein-Stein model
$=$ $V_{t}+\rho\nu\int_{t}^{T}e^{-\kappa(s-t)}\,\mathrm{d}W_{s}-\rho\nu\int_{t}^{T}V_{s}\mathrm{e}^{-\kappa(s-t)}\,\mathrm{d}s$
@@
From this last equation we also deduce that the second order Malliavin derivative of $G(t,T)$ vanishes.